# references.bib

### v1

@@ 5.3.1. Proactive Approaches
By predicting the resource requirements of workloads (jeatsa_casy_2022; stewart_performance_2022), or by observing workload characteristics to identify beneficial static configurations (noll_accelerating_2018), proactive approaches improve system goals ahead of time.
CASY (jeatsa_casy_2022) by jeatsa_casy_2022 is a cache allocation controller for function-as-a-service (FaaS) systems that uses machine learning to build a cache usage profile of functions with regard to their input values. This allows the controller to tune functions’ allocation of cache by examining the incoming input and thus improve system-wide performance.
With the use of deep learning, stewart_performance_2022 (stewart_performance_2022) profile web services’ cache usage characteristics. This allows them to propose their short-term cache allocation controller that observes response times of web services and temporarily allocates additional LLC to particularly slow requests in order to minimize SLO infringement.
Database management systems (DBMS) are required to handle large amounts of queries and data, but all operations are not equal in cache sensitivity and efficiency as noted by noll_accelerating_2018 (noll_accelerating_2018). They find that certain memory-intensive operations such as a column scan are insensitive to available cache, while simultaneously thrashing existing cache contents. Thus, they propose allocating the minimum possible cache to worker threads that handle such queries which reduces overall cache pollution and thus increases system performance by, in their evaluation, up to $38\%$.
@@ 5.3.2. Reactive Approaches
Through observation of both application metrics and system metrics, DICER (nikas_dicer_2019) classifies workloads based on their sensitivity to LLC availability. Additionally, it monitors MB saturation and detects workload phase changes by identifying changes in LLC utilization. The controller proposed by nikas_dicer_2019 uses the gathered data to heuristically increase system utilization by improving BE job performance without allowing LC workloads to break QoS targets.
@@
Virtual network functions (VNFs) often have strict performance guarantees to meet. chintapalli_restrain_2022 (chintapalli_restrain_2022) argue that to be able to co-locate multiple VNFs with other services while ensuring QoS, isolation of resources is crucial. They present RESTRAIN, a dynamic LLC and MB allocation scheme that continuously classifies VNFs into one of four categories using a modified technique borrowed from park_copart_2019 (park_copart_2019). The categories used are “donor”, “receiver”, “donor and receiver”, and “none”. A donor is a workload that is able to release a certain resource and still meet QoS, while receivers need more resources to meet QoS. Donors and receivers are workloads that require more of one resource while being in excess of another. Workloads that are at a resource allocation are classified as none. The controller builds upon and is compared to ResQ (tootoonchian_resq_2018).
@@
With large scale data-centers in mind, duan_improving_2021 (duan_improving_2021) propose a machine learning-based workload classification model. This model is later used by a local search server re-consolidation algorithm. The controller utilizes LLC partitioning and core allocation to guarantee workload performance during re-consolidation from overloaded machines. This work is continued and extended by the same authors in duan_improving_2022 (duan_improving_2022).
@@
With the goal of increasing system utilization by co-locating a greater number of BE jobs without breaking LC service SLOs, patel_clite_2020 (patel_clite_2020) propose CLITE. CLITE is a memory bandwidth and LLC partitioning controller that uses Bayesian optimization to find an optimal configuration of the system. In their case, the controller maximizes the performance of BE jobs while LC applications still meet their QoS requirements.
@@
Identifying the most critical resource of a workload at runtime can be difficult. pons_effect_2022 (pons_effect_2022) try to show how different types of workloads behave when constrained by different kinds of resources. This gives an insight into how resource allocation features behave when used in conjunction. For example, they show that a reduction in LLC availability may increase the memory bandwidth used by a process. This allows them to present a number of key takeaways and guidelines that may help guide cloud operators to make more efficient use of their hardware by reducing resource allocation to insensitive workloads.
@@
gureya_generalizing_2021 (gureya_generalizing_2021) recognize that Intel’s MBA is limited to single socket control only due to the details of its implementation. In an attempt to bring holistic memory bandwidth allocation to multi-socket systems, they propose BALM. BALM is a controller that combines a cross-socket page migration scheme with knowledge of workloads’ SLO targets. This allows the controller to achieve better overall system performance through memory bandwidth partitioning and socket migration of workloads while maintaining SLO targets.
@@
By allocating an exclusive cache partition to a high-priority workload while other applications share the rest, the QoS of the high-priority workload can be maintained. This is what papadakis_improving_2017 propose with their heuristic Dynamic Cache-Partitioning QoS scheme (papadakis_improving_2017). The controller observes the performance of the high-priority workload and adjusts its LLC allocation such that QoS is maintained while maximizing LLC available to remaining system workloads.
@@
qin_locpart_2017 approach LLC partitioning by allocating each workload a separate exclusive piece of LLC in Locpart (qin_locpart_2017). By monitoring the QoS of LC workloads, LLC may be redistributed to or from LC workloads to BE jobs in such a way that LC services meet their performance targets and BE job performance is maximized. A limitation of the approach is that the number of available COSs dictates the maximum number of total workloads assigned to a system.
@@ 5.3.3. Combined Approaches
With the goal of increasing system performance while maintaining LC workload SLO requirements, qiu_slo_2021 (qiu_slo_2021) propose CoCo. CoCo is an SLO-aware LLC and memory bandwidth partitioning controller that profiles workloads’ sensitivity to such scarce resources. With this knowledge, it dynamically allocates resources based on available SLO overhead to increase overall system utilization and performance.
@@
nejat_coordinated_2020 have total system efficiency in mind when proposing their resource management controller (nejat_coordinated_2020). By observing the SLOs and metrics of workloads, their controller is able to reduce power consumption while still meeting those QoS requirements. This is achieved by modeling the cache and power usage behavior of workloads and then partitioning LLC and scaling core voltages to reduce energy consumption while still allowing applications to meet their intended targets.
@@
By gathering multiple co-located parts of the same LC service into a new abstraction called a Servpod, zhao_rhythm_2020 (zhao_rhythm_2020) are able to improve system utilization while ensuring that LC workloads meet their SLOs. This is achieved by offline profiling and classification of Servpods based on their contributions to overall system tail-latency and sensitivity to shortage of shared scarce resources. They call their scheme Rhythm, which partitions LLC, manages core voltages, etc., with the goal of increasing utilization without LC SLO breakage.
@@
By using the result from offline profiling of VNFs prior to deployment, ResQ (tootoonchian_resq_2018) is able to make greedy decisions regarding LLC allocation optimizations. The controller, proposed by tootoonchian_resq_2018, ensures VNF QoS requirements are met while maximizing the performance of co-located BE workloads.
@@
To allow workload owners to specify SLO targets instead of resource allocation requests in the cloud, metsch_intent-driven_2023 (metsch_intent-driven_2023) propose their intent driven orchestrator. The system learns from application behavior and how changes in the allocation of resources such as LLC affect SLO metrics. This allows the controller to predict how changes in resource allocation will affect application performance and make resource partitioning decisions accordingly.
@@
Due to strict SLO requirements, VNFs are often placed on underutilized machines so that their performance can be guaranteed. li_rldrm_2020 (li_rldrm_2020) propose RLDRM in an attempt to improve utilization of such machines. RLDRM is a closed-loop deep reinforcement learning-based LLC partitioning controller. By learning from the behavior of workloads the controller can make sure the high-priority VNFs have enough LLC ways to meet their SLO requirements while allowing a greater number of BE jobs to consume the remaining resources.
@@
After identifying the shortcomings of Intel’s MBA, zhang_libra_2021 (zhang_libra_2021) propose a custom memory bandwidth partitioning mechanism they call LIBRA. By custom manipulation of the read pending queue, finer-grained control is achieved than what is available through MBA. LIBRA still leverages CAT as a mechanism of LLC partitioning.
@@
Through prior profiling of VNF sensitivity to cache and memory bandwidth constraints, chintapalli_ravin_2023 (chintapalli_ravin_2023) propose RAVIN. The heuristic VNF orchestrator schedules VNFs across machines while considering available cache and memory bandwidth such that throughput guarantees may be met. This is done using the previously gathered profiling information which provides the orchestrator with minimum resource requirements depending on desired throughput guarantees. This allows RAVIN to efficiently utilize available hardware through more efficient packing of VNFs.
@@
To ensure that the worst-case execution time (WCET) of real-time workloads stays within requirements, xu_holistic_2019 propose CaM (xu_holistic_2019). The three-phase heuristic algorithm groups workloads based on their sensitivity to resource shortage before it load-balances them across the cores of a system. LLC and memory BW (using MemGuard (yun_memguard_2013)) are then allocated such that WCET targets are fulfilled.
@@ 5.3.4. Performance Slack Management Approaches
With knowledge of an LC workload’s QoS requirements and prior offline profiling, Dirigent (zhu_dirigent_2016) may manage LLC in such a way that co-located BE jobs’ performance can be maximized and thus increase overall system utilization. zhu_dirigent_2016 approach the problem by predicting the remaining runtime of workloads using data gathered during the prior profiling. By comparing how it correlates to set deadlines, resources may be adjusted dynamically so that the performance of LC workloads meets set targets while allowing an increased amount of resources to co-located jobs.
With the premise that only a single LC workload is located on any one machine to guarantee QoS, chen_parties_2019 (chen_parties_2019) propose PARTIES. PARTIES is a QoS-aware resource partitioning controller that enables the co-location and QoS enforcement of multiple LC workloads while maximizing the performance of BE jobs to improve system utilization. This is achieved by observing the performance slack of LC workloads and increasing resource allocation of services about to break their targets.
@@
In one of the first works utilizing platform features, lo_heracles_2015 propose Heracles (lo_heracles_2015), a resource partitioning controller that monitors QoS metrics of an LC workload. By observing that LC workload’s performance slack, its resource allocations may be adjusted such that just enough resources are available to meet its SLO targets. This allows the remaining resources to be distributed among BE jobs, thus improving system utilization and the performance of those services.
@@
By being aware of workload QoS targets, a controller may locate LLC resources in a way that decreases the overall power consumption of a system through the management of performance slack. This is what nejat_cooperative_2022 (nejat_cooperative_2022) propose with their Cooperative Slack Management scheme. It identifies opportunities to generate slack with minimal use of energy. This slack may later be consumed at times of less power-efficient operation and thus increase energy efficiency over time.
@@
Sturgeon (pang_adaptive_2021) is a multi-level resource management system that manages BE and LC workload placement at the cluster level in addition to providing machine-level resource management with a focus on achieving higher performance while not increasing energy consumption. They propose using machine learning models to predict both the performance and power consumption of heuristic changes to resource partitioning based on performance slack. This is to ensure services stay within SLO requirements and that the system does not breach power usage constraints.
@@
Avalon (chen_avalon_2019) is a resource management scheme proposed by chen_avalon_2019 that attempts to improve system utilization by providing “just enough” resources to LC services while ensuring QoS targets are met. This is done by predicting the resource requirements for each incoming LC workload query using an ML model and allocating resources appropriately. The execution time slack of queries is monitored and heuristic adjustments to resource allocation are made if necessary to meet QoS targets. This improves utilization by providing co-located BE jobs with increased resources while ensuring LC QoS.
@@ 5.3.5. Summary
Being application-aware means a higher configuration overhead for controllers managing a heterogeneous set of workloads. However, it does allow controllers to optimize the resource partitioning of a system while maintaining workload SLO targets. Publications covered in this section have seen improvements in LC workload throughput by $31.7\%$ (zhao_rhythm_2020), system utilization by $28.9\%$ (chen_avalon_2019), and serverless function execution time by $11\%$ (jeatsa_casy_2022). It is therefore clear that application-aware management of QoS enforcement features may greatly benefit execution environments not already configured to limit congestion over shared resources.
The problem of shared resource management is approached in the context of virtual network functions by (li_rldrm_2020; chintapalli_restrain_2022; tootoonchian_resq_2018; chintapalli_ravin_2023). It is common that such functions are constrained by strict QoS requirements, and are thus traditionally deployed on machines with relatively low total utilization (li_rldrm_2020). By guaranteeing scarce resource access, utilization of such machines may be greatly improved for reduced cost of ownership and increased efficiency.
By observing LC workload deadlines and by efficiently managing performance slack, efficiency (nejat_cooperative_2022; pang_adaptive_2021), server utilization (zhu_dirigent_2016; chen_parties_2019; chen_avalon_2019), or system performance (lo_heracles_2015) may be improved. This is done by identifying times at which slack may be increased at a low cost such that it may be consumed at a later time when the cost is higher.

### v2

@@ 5.3.1. Proactive Approaches
By predicting the resource requirements of workloads (jeatsa_casy_2022, 51, 97), or by observing workload characteristics to identify beneficial static configurations (noll_accelerating_2018, 71), proactive approaches improve system goals ahead of time.
CASY (jeatsa_casy_2022, 51) by (jeatsa_casy_2022, 51) is a cache allocation controller for function-as-a-service (FaaS) systems that uses machine learning to build a cache usage profile of functions with regard to their input values. This allows the controller to tune functions’ allocation of cache by examining the incoming input and thus improve system-wide performance.
With the use of deep learning, (stewart_performance_2022, 97) (stewart_performance_2022, 97) profile web services’ cache usage characteristics. This allows them to propose their short-term cache allocation controller that observes response times of web services and temporarily allocates additional LLC to particularly slow requests in order to minimize SLO infringement.
Database management systems (DBMS) are required to handle large amounts of queries and data, but all operations are not equal in cache sensitivity and efficiency as noted by (noll_accelerating_2018, 71) (noll_accelerating_2018, 71). They find that certain memory-intensive operations such as a column scan are insensitive to available cache, while simultaneously thrashing existing cache contents. Thus, they propose allocating the minimum possible cache to worker threads that handle such queries which reduces overall cache pollution and thus increases system performance by, in their evaluation, up to $38\%$.
@@ 5.3.2. Reactive Approaches
Through observation of both application metrics and system metrics, DICER (nikas_dicer_2019, 70) classifies workloads based on their sensitivity to LLC availability. Additionally, it monitors MB saturation and detects workload phase changes by identifying changes in LLC utilization. The controller proposed by (nikas_dicer_2019, 70) uses the gathered data to heuristically increase system utilization by improving BE job performance without allowing LC workloads to break QoS targets.
@@
Virtual network functions (VNFs) often have strict performance guarantees to meet. (chintapalli_restrain_2022, 18) (chintapalli_restrain_2022, 18) argue that to be able to co-locate multiple VNFs with other services while ensuring QoS, isolation of resources is crucial. They present RESTRAIN, a dynamic LLC and MB allocation scheme that continuously classifies VNFs into one of four categories using a modified technique borrowed from (park_copart_2019, 74) (park_copart_2019, 74). The categories used are “donor”, “receiver”, “donor and receiver”, and “none”. A donor is a workload that is able to release a certain resource and still meet QoS, while receivers need more resources to meet QoS. Donors and receivers are workloads that require more of one resource while being in excess of another. Workloads that are at a resource allocation are classified as none. The controller builds upon and is compared to ResQ (tootoonchian_resq_2018, 102).
@@
With large scale data-centers in mind, (duan_improving_2021, 24) (duan_improving_2021, 24) propose a machine learning-based workload classification model. This model is later used by a local search server re-consolidation algorithm. The controller utilizes LLC partitioning and core allocation to guarantee workload performance during re-consolidation from overloaded machines. This work is continued and extended by the same authors in (duan_improving_2022, 25) (duan_improving_2022, 25).
@@
With the goal of increasing system utilization by co-locating a greater number of BE jobs without breaking LC service SLOs, (patel_clite_2020, 76) (patel_clite_2020, 76) propose CLITE. CLITE is a memory bandwidth and LLC partitioning controller that uses Bayesian optimization to find an optimal configuration of the system. In their case, the controller maximizes the performance of BE jobs while LC applications still meet their QoS requirements.
@@
Identifying the most critical resource of a workload at runtime can be difficult. (pons_effect_2022, 81) (pons_effect_2022, 81) try to show how different types of workloads behave when constrained by different kinds of resources. This gives an insight into how resource allocation features behave when used in conjunction. For example, they show that a reduction in LLC availability may increase the memory bandwidth used by a process. This allows them to present a number of key takeaways and guidelines that may help guide cloud operators to make more efficient use of their hardware by reducing resource allocation to insensitive workloads.
@@
(gureya_generalizing_2021, 36) (gureya_generalizing_2021, 36) recognize that Intel’s MBA is limited to single socket control only due to the details of its implementation. In an attempt to bring holistic memory bandwidth allocation to multi-socket systems, they propose BALM. BALM is a controller that combines a cross-socket page migration scheme with knowledge of workloads’ SLO targets. This allows the controller to achieve better overall system performance through memory bandwidth partitioning and socket migration of workloads while maintaining SLO targets.
@@
By allocating an exclusive cache partition to a high-priority workload while other applications share the rest, the QoS of the high-priority workload can be maintained. This is what (papadakis_improving_2017, 73) propose with their heuristic Dynamic Cache-Partitioning QoS scheme (papadakis_improving_2017, 73). The controller observes the performance of the high-priority workload and adjusts its LLC allocation such that QoS is maintained while maximizing LLC available to remaining system workloads.
@@
(qin_locpart_2017, 83) approach LLC partitioning by allocating each workload a separate exclusive piece of LLC in Locpart (qin_locpart_2017, 83). By monitoring the QoS of LC workloads, LLC may be redistributed to or from LC workloads to BE jobs in such a way that LC services meet their performance targets and BE job performance is maximized. A limitation of the approach is that the number of available COSs dictates the maximum number of total workloads assigned to a system.
@@ 5.3.3. Combined Approaches
With the goal of increasing system performance while maintaining LC workload SLO requirements, (qiu_slo_2021, 85) (qiu_slo_2021, 85) propose CoCo. CoCo is an SLO-aware LLC and memory bandwidth partitioning controller that profiles workloads’ sensitivity to such scarce resources. With this knowledge, it dynamically allocates resources based on available SLO overhead to increase overall system utilization and performance.
@@
(nejat_coordinated_2020, 69) have total system efficiency in mind when proposing their resource management controller (nejat_coordinated_2020, 69). By observing the SLOs and metrics of workloads, their controller is able to reduce power consumption while still meeting those QoS requirements. This is achieved by modeling the cache and power usage behavior of workloads and then partitioning LLC and scaling core voltages to reduce energy consumption while still allowing applications to meet their intended targets.
@@
By gathering multiple co-located parts of the same LC service into a new abstraction called a Servpod, (zhao_rhythm_2020, 123) (zhao_rhythm_2020, 123) are able to improve system utilization while ensuring that LC workloads meet their SLOs. This is achieved by offline profiling and classification of Servpods based on their contributions to overall system tail-latency and sensitivity to shortage of shared scarce resources. They call their scheme Rhythm, which partitions LLC, manages core voltages, etc., with the goal of increasing utilization without LC SLO breakage.
@@
By using the result from offline profiling of VNFs prior to deployment, ResQ (tootoonchian_resq_2018, 102) is able to make greedy decisions regarding LLC allocation optimizations. The controller, proposed by (tootoonchian_resq_2018, 102), ensures VNF QoS requirements are met while maximizing the performance of co-located BE workloads.
@@
To allow workload owners to specify SLO targets instead of resource allocation requests in the cloud, (metsch_intent-driven_2023, 64) (metsch_intent-driven_2023, 64) propose their intent driven orchestrator. The system learns from application behavior and how changes in the allocation of resources such as LLC affect SLO metrics. This allows the controller to predict how changes in resource allocation will affect application performance and make resource partitioning decisions accordingly.
@@
Due to strict SLO requirements, VNFs are often placed on underutilized machines so that their performance can be guaranteed. (li_rldrm_2020, 57) (li_rldrm_2020, 57) propose RLDRM in an attempt to improve utilization of such machines. RLDRM is a closed-loop deep reinforcement learning-based LLC partitioning controller. By learning from the behavior of workloads the controller can make sure the high-priority VNFs have enough LLC ways to meet their SLO requirements while allowing a greater number of BE jobs to consume the remaining resources.
@@
After identifying the shortcomings of Intel’s MBA, (zhang_libra_2021, 122) (zhang_libra_2021, 122) propose a custom memory bandwidth partitioning mechanism they call LIBRA. By custom manipulation of the read pending queue, finer-grained control is achieved than what is available through MBA. LIBRA still leverages CAT as a mechanism of LLC partitioning.
@@
Through prior profiling of VNF sensitivity to cache and memory bandwidth constraints, (chintapalli_ravin_2023, 19) (chintapalli_ravin_2023, 19) propose RAVIN. The heuristic VNF orchestrator schedules VNFs across machines while considering available cache and memory bandwidth such that throughput guarantees may be met. This is done using the previously gathered profiling information which provides the orchestrator with minimum resource requirements depending on desired throughput guarantees. This allows RAVIN to efficiently utilize available hardware through more efficient packing of VNFs.
@@
To ensure that the worst-case execution time (WCET) of real-time workloads stays within requirements, (xu_holistic_2019, 112) propose CaM (xu_holistic_2019, 112). The three-phase heuristic algorithm groups workloads based on their sensitivity to resource shortage before it load-balances them across the cores of a system. LLC and memory BW (using MemGuard (yun_memguard_2013, 118)) are then allocated such that WCET targets are fulfilled.
@@ 5.3.4. Performance Slack Management Approaches
With knowledge of an LC workload’s QoS requirements and prior offline profiling, Dirigent (zhu_dirigent_2016, 125) may manage LLC in such a way that co-located BE jobs’ performance can be maximized and thus increase overall system utilization. (zhu_dirigent_2016, 125) approach the problem by predicting the remaining runtime of workloads using data gathered during the prior profiling. By comparing how it correlates to set deadlines, resources may be adjusted dynamically so that the performance of LC workloads meets set targets while allowing an increased amount of resources to co-located jobs.
@@
In one of the first works utilizing platform features, (lo_heracles_2015, 61) propose Heracles (lo_heracles_2015, 61), a resource partitioning controller that monitors QoS metrics of an LC workload. By observing that LC workload’s performance slack, its resource allocations may be adjusted such that just enough resources are available to meet its SLO targets. This allows the remaining resources to be distributed among BE jobs, thus improving system utilization and the performance of those services.
@@
By being aware of workload QoS targets, a controller may locate LLC resources in a way that decreases the overall power consumption of a system through the management of performance slack. This is what (nejat_cooperative_2022, 68) (nejat_cooperative_2022, 68) propose with their Cooperative Slack Management scheme. It identifies opportunities to generate slack with minimal use of energy. This slack may later be consumed at times of less power-efficient operation and thus increase energy efficiency over time.
@@
Sturgeon (pang_adaptive_2021, 72) is a multi-level resource management system that manages BE and LC workload placement at the cluster level in addition to providing machine-level resource management with a focus on achieving higher performance while not increasing energy consumption. They propose using machine learning models to predict both the performance and power consumption of heuristic changes to resource partitioning based on performance slack. This is to ensure services stay within SLO requirements and that the system does not breach power usage constraints.
@@
Avalon (chen_avalon_2019, 14) is a resource management scheme proposed by (chen_avalon_2019, 14) that attempts to improve system utilization by providing “just enough” resources to LC services while ensuring QoS targets are met. This is done by predicting the resource requirements for each incoming LC workload query using an ML model and allocating resources appropriately. The execution time slack of queries is monitored and heuristic adjustments to resource allocation are made if necessary to meet QoS targets. This improves utilization by providing co-located BE jobs with increased resources while ensuring LC QoS.
@@ 5.3.5. Summary
Being application-aware means a higher configuration overhead for controllers managing a heterogeneous set of workloads. However, it does allow controllers to optimize the resource partitioning of a system while maintaining workload SLO targets. Publications covered in this section have seen improvements in LC workload throughput by $31.7\%$ (zhao_rhythm_2020, 123), system utilization by $28.9\%$ (chen_avalon_2019, 14), and serverless function execution time by $11\%$ (jeatsa_casy_2022, 51). It is therefore clear that application-aware management of QoS enforcement features may greatly benefit execution environments not already configured to limit congestion over shared resources.
The problem of shared resource management is approached in the context of virtual network functions by (li_rldrm_2020, 57, 18, 102, 19). It is common that such functions are constrained by strict QoS requirements, and are thus traditionally deployed on machines with relatively low total utilization (li_rldrm_2020, 57). By guaranteeing scarce resource access, utilization of such machines may be greatly improved for reduced cost of ownership and increased efficiency.
By observing LC workload deadlines and by efficiently managing performance slack, efficiency (nejat_cooperative_2022, 68, 72), server utilization (zhu_dirigent_2016, 125, 17, 14), or system performance (lo_heracles_2015, 61) may be improved. This is done by identifying times at which slack may be increased at a low cost such that it may be consumed at a later time when the cost is higher.